# Generalized Short Path Algorithms: Towards Super-Quadratic Speedup over Markov Chain Search for Combinatorial Optimization

### v1

@@ 4.2.2 The Long Jump
As should be apparent from Theorem 4.3, at a constant $b$ the existence of a speedup over Grover is determined solely by $\frac{\Delta_{P}}{\lvert E^{*}\rvert}$. Clearly, we must have that $\frac{\Delta_{P}}{\lvert E^{*}\rvert}=\Omega\left(\frac{1}{\ln(1/\pi(E^{*}))}\right)$, at least for a problem with at least exponential runtime. If this does not hold, then the runtime goes to zero asymptotically, an absurdity. However, it has not been directly shown that the derived runtime cannot lead to this contradiction. To put one’s mind at ease, we present the following result.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain. Suppose $b^{*}$ and $\gamma$ in Theorem 4.3 are constants, and $\lVert H\rVert_{2}=\lvert E^{*}\rvert$. If
[…]
@@
Assume that $\Delta$ is a global upper bound, such that for all $x\in\mathcal{S}(n)$, $\mathbb{E}_{y\sim_{P}x}[H(y)]\leq H(x)+\Delta$. It is easy to observe that for any random variable $X$ taking values in $\mathcal{S}(n)$, it holds that $\mathbb{E}_{Y\sim_{P}X}[H(Y)]\leq\mathbb{E}_{X}[H(X)]+\Delta$. Now let $x_{\ast}$ be a global minimum of $H$ (with corresponding energy $E^{\ast}$) and consider taking $T=\lceil M(n)\rceil$ steps of $P$ starting from $x^{\ast}$, with the random state after $t\in[1,T]$ steps being denoted $X_{t}$. By induction, it is easy to see that $E_{X_{T}}[H(X_{T})]\leq E^{\ast}+\Delta T$. From the definition of mixing time however, it follows that $\mathbb{E}_{X_{T}}[H(X_{T})]\geq\mathbb{E}_{\pi}[H(x)]-\frac{\lvert E^{\ast}\rvert}{100}$. Denoting $\mathbb{E}_{\pi}[H(x)]$ by $\bar{E}$ it follows that $\Delta\geq\frac{\bar{E}-E^{\ast}}{T}-\frac{\lvert E^{*}\rvert}{100T}=\Omega\left(\frac{\lvert E^{*}\rvert}{M(n)}\right)$. So $\frac{\tilde{\Delta}_{P}}{\lvert E^{\ast}\rvert}=\Omega\left(\frac{1}{M(n)}\right)$.
@@
Note that in [DPCB23], the authors state an additional technical condition that is of course easy to satisfy, which is $\mathbb{E}_{\pi}[H(x)]=0$. The main reason for setting $\mathbb{E}_{\pi}[H(x)]=0$ is to ensure that $E^{\ast}<0$ and for the ease of proving the tail bounds. In our setting, for arbitrary $\pi$, the expectation may need to be estimated if used as a shift. However, the shift is not necessary to run the algorithm if the $E^{\ast}<0$ condition is already satisfied. The mean just appears as component of the runtime. Also, Corollaries 4.9 and 4.10 for the tail bounds do not assume this shift.
@@ 5.1 Optimization with Fixed Hamming Weight: Transposition Walk
The $k$-particle Bernoulli-Laplace diffusion or Transposition Walk on $n$ sites is a random walk on the space of Hamming weight $k$ bistrings. For our case we will work with ±$\pm 1$ strings or “spin configurations” $x$ and define the Hamming weight $\lvert x\rvert$ as the number of $+1$’s. Formally, $\mathcal{X}=\{x\in\{-1,1\}^{n}:\lvert x\rvert=k\}$. A single step of BL consists of choosing, uniformly at random, a transposition that swaps some $x_{j}=1$ with another $x_{i}=-1$.
@@
where $X_{j},Y_{j}$ denote the Pauli operators applied to qubit $j$. This is commonly called the complete-graph $XY$ mixer. Note that there is equality between $D(P)$ and $P$ because the walk is symmetric. The ground state of $-D(P)$ is the uniform superposition over Hamming-weight $k$ computational basis states, and thus encodes the stationary distribution of the transposition walk. Note that $|\sqrt{\pi}\rangle$ is just the Hamming-weight $k$ Dicke state, which can be prepared efficiently [BE19].
@@
$\mathcal{D}(\psi,\psi)\geq\frac{n}{k(n-k)\tau_{LS}}\mathrm{Ent}(\psi^{2}).$
There is also a universal constant $\tau_{0}$ such that
$\tau_{LS}\leq\tau_{0}\log\left(\frac{n}{\min(k,n-k)}\right).$
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the $k$-particle transposition walk on $n$ sites. Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy $E^{\ast}$. If $\lVert H\rVert_{P}=\mathcal{O}(1)$, and
@@
then there exists a short-path algorithm with runtime
@@
We apply the above result to a Hamming-weight constrained version of MaxCut over Erdős-Rényi graphs, which we call MaxCut-Hamming. One well-known special case is Hamming-weight $\frac{n}{2}$ called MaxBisection.
5.1.1 Hamming-weight Constrained MaxCut
Consider a graph $G(\mathcal{N},\mathcal{E})$ with vertex set $\mathcal{N}:=[n]$ and edge set $\mathcal{E}$. We assume $G$ is drawn from the Erdős-Rényi ensemble $\mathcal{G}\left(n,\frac{p}{n-1}\right)$ for a constant $p$, i.e., each edge $e_{ij}\in\mathcal{E}$ for $(i,j)\in[n]\times[n]$ is created with probability $\frac{p}{n-1}$ such that $G$ has an average degree of $p$. We are interested in solving the Maximum Bisection problem:
@@ 5.1.1 Hamming-weight Constrained MaxCut
which gives that $\frac{\lvert E^{*}\rvert}{\Delta_{P}}=\mathcal{O}(k)$. However, $\ln\binom{n}{k}=\Theta(k\ln(n/k))$ for $k=o(n)$. Assuming a tail bound, this leads to a speedup that is falling with $n$. Specifically, the speedup is falling with $\frac{1}{\ln(n)}$, which is reminiscent of the running time achieved by [Has18a]. We summarize the two cases in the following theorem
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the $k$-particle transposition walk on $n$ sites. Let $H:\left\{-1,1\right\}^{n}\mapsto\mathbb{R}$ be a diagonal Hamiltonian encoding the MaxCut-Hamming cost function for a graph $G\sim\mathcal{G}\left(n,\frac{p}{n-1}\right)$. Then either of the following runtimes hold depending on $k\leq n$.
  • If $k=\Theta(n)$, then there exists a short-path algorithm with runtime
@@
  • if $k=o(n)$, then there exists a short-path algorithm with runtime
@@ 5.2 Glauber Dynamics
Glauber dynamics is a Markov Chain algorithm designed to sample from the Gibbs distribution of a system, particularly in spin models like the Ising or hard-core model [Gla63]. The Gibbs measure $\pi$ for a system with configuration space $\mathcal{X}$ and Hamiltonian $H(x)$ is defined as
$\pi(x)=\frac{\exp(-\beta H(x))}{Z(\beta)}$
@@
Glauber dynamics generates a Markov chain with this measure as its stationary distribution by sequentially updating a single site (spin or vertex) according marginal distribution. Specifically, the update proceeds as follows: (i) a site (vertex) $v$ is chosen uniformly at random; (ii) the states of all other sites $u\neq v$ remain unchanged; (iii) the new state of $v$ is sampled from the marginal distribution of $v$ conditioned on its neighbors.
@@
The efficiency of Gibbs sampling in this case is related to how fast the Glauber dynamics mixes to its stationary distribution. In fact, approximate Gibbs sampling is tightly connected to partition function estimation in terms of computational complexity [ŠVV09], both of which are closely related to statistical phase transitions. These transitions correspond to the uniqueness/non-uniqueness threshold on an infinite $d$-regular tree which captures whether the root of the tree is affected by the leaves. In the uniqueness regime, correlations decay rapidly, allowing efficient approximation of the partition function. However, beyond the non-uniqueness threshold, long range correlations emerge and no polynomial-time algorithm can approximate the partition function.
@@
In particular, for the hardcore model with fugacity parameter $\lambda=e^{\beta}$, [Wei06] presented a fully polynomial time approximation scheme ($\mathsf{FPTAS}$) for computing the partition function a graphs with maximum degree $d$ when $\lambda\leq(1-\delta)\lambda_{c}$ where $\lambda_{c}=\frac{(d-1)^{(d-1)}}{(d-2)^{d}}$ is the corresponding uniqueness threshold. On the other hand,[Sly10] proved that is no fully polynomial randomized approximation scheme ($\mathsf{FPRAS}$) to approximate the partition function when $\lambda>\lambda_{c}$ unless $\mathsf{NP}=\mathsf{RP}$ confirming the main conjecture of [MWW09]. Similarly, for Ising model the phase transition occurs at $\beta_{c}=\frac{d-2}{d}$ for the antiferromagnetic case and $\beta_{c}=\frac{d}{d-2}$ for ferromagnetic case.
We first present our main result for the Glauber dynamics and then show that the necessary conditions hold for hardcore-model (maximum independent set problem) and Ising model.
Theorem 5.7.
Let $P$ be a Glauber dynamics chain on $\mathcal{X}$ that satisfies an $\omega$ log-Sobolev inequality with stationary distribution $\pi$, and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
$E^{\ast}\coloneqq\min_{x\in\mathcal{X}}H(x).$
If $\|H\|_{P}=\mathcal{O}(1)$ and $\omega^{-1},\lvert E^{*}\rvert,$ and $\ln(1/\pi(E^{*}))$ are all $\Theta(n)$, then the short-path algorithm applied to the Glauber chain has a super-quadratic speedup over Markov Chain search with Glauber dynamics.
Proof.
This directly follows from application of Theorem 4.3. ∎
@@
Let $P$ be a Glauber dynamics chain on $\mathcal{X}$ and let $H:\mathcal{X}\mapsto\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
@@
If the following holds for all $x,x^{\prime}$ such that $P(x,x^{\prime})>0$
$\lvert H(x^{\prime})-H(x)\rvert=\mathcal{O}(1),$ (41)
then $\|H\|_{P}=\mathcal{O}(1)$.
@@
$\|H\|_{P}$ $=\max_{x}\sum_{x^{\prime}}P(x,x^{\prime})(H(x^{\prime})-H(x))^{2}\leq\max_{x,x^{\prime}}\lvert H(x^{\prime})-H(x)\rvert^{2}=\mathcal{O}(1).$
5.2.1 Maximum Independent Set Problem

### v2

@@ 4.2.2 The Long Jump
As mentioned in Corollary 4.5, at a constant $b$ the existence of an asymptotic speedup over Grover is determined solely by $\frac{\Delta_{P}}{\lvert E^{*}\rvert}$. Clearly, we must have that $\frac{\Delta_{P}}{\lvert E^{*}\rvert}=\Omega\left(\frac{1}{\ln(1/\pi(E^{*}))}\right)$, at least for a problem with at least exponential runtime. If this does not hold, then the runtime goes to zero asymptotically, an absurdity. However, it has not been directly shown that the derived runtime cannot lead to this contradiction. To put one’s mind at ease, we present the following result.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a Markov chain with log-Sobolev constant lower bounded by $\omega$. Suppose $b^{*}$ and $\gamma$ in Theorem 4.2 are $\Theta(1)$, and $\lVert H\rVert_{2}=\lvert E^{*}\rvert$. If
$\mathbb{E}_{y\underset{P}{\sim}x}[H(y)]\leq H(x)+\tilde{\Delta}_{P},\quad\forall x\in\mathcal{X},$ (33)
@@
Assume that $\tilde{\Delta}_{P}$ is a global upper bound, such that for all $x\in\mathcal{S}(n)$, $\mathbb{E}_{y\sim_{P}x}[H(y)]\leq H(x)+\tilde{\Delta}_{P}$. It is easy to observe that for any random variable $X$ taking values in $\mathcal{S}(n)$, it holds that $\mathbb{E}_{Y\sim_{P}X}[H(Y)]\leq\mathbb{E}_{X}[H(X)]+\tilde{\Delta}_{P}$. Now let $x_{\ast}$ be a global minimum of $H$ (with corresponding energy $E^{\ast}$) and consider taking $T=\lceil M(n)\rceil$ steps of $P$ starting from $x^{\ast}$, with the random state after $t\in[1,T]$ steps being denoted $X_{t}$. By induction, it is easy to see that $E_{X_{T}}[H(X_{T})]\leq E^{\ast}+\tilde{\Delta}_{P}T$. From the definition of mixing time however, it follows that $\mathbb{E}_{X_{T}}[H(X_{T})]\geq\mathbb{E}_{\pi}[H(x)]-\frac{\lvert E^{\ast}\rvert}{100}$. Denoting $\mathbb{E}_{\pi}[H(x)]$ by $\bar{E}$ it follows that $\tilde{\Delta}_{P}\geq\frac{\bar{E}-E^{\ast}}{T}-\frac{\lvert E^{*}\rvert}{100T}=\Omega\left(\frac{\lvert E^{*}\rvert}{M(n)}\right)$. So $\frac{\tilde{\Delta}_{P}}{\lvert E^{\ast}\rvert}=\Omega\left(\frac{1}{M(n)}\right)$.
@@
Note that in [DPCB23], the authors state an additional technical condition that is of course easy to satisfy, which is $\mathbb{E}_{\pi}[H(x)]=0$. The main reason for setting $\mathbb{E}_{\pi}[H(x)]=0$ is to ensure that $E^{\ast}<0$ and for the ease of proving the tail bounds. In our setting, for arbitrary $\pi$, the expectation may need to be estimated if used as a shift. However, the shift is not necessary to run the algorithm if the $E^{\ast}<0$ condition is already satisfied. The mean just appears as component of the runtime. Also, Corollaries 4.8 and 4.9 for the tail bounds do not assume this shift.
Lastly, we obtain the following self-evident corollary of Theorem 4.20 when applied to Equation (9)
Corollary 4.22.
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be a reversible, aperiodic Markov chain. Let $\nu$ be a lower bound on either the log-Sobolev or Poincaré constant of $P$. Given positive parameters $\eta<1$, $b<1$, $\alpha<(1-b)/2$. Suppose that $(H,g_{\eta})$ has the $\alpha$-subdepolarizing property, and that there exists an integer $\ell$ such that
$\max\left(\frac{4\ln(V)}{\nu},\frac{3}{\alpha^{2}}\right)<\ell\leq\frac{\langle\psi_{b}|\sqrt{\pi}\rangle}{b\sqrt{\mathbb{P}_{\pi}(E\leq(1-\eta)E^{*})}}.$
If $|\psi_{b}\rangle$ is the unique ground state of $H_{b}$ and the spectral gap of $H_{b}$ is lower bounded by $\frac{\nu}{2}$, then the runtime of the short-path algorithm is bounded by
$\mathcal{O}\left(\textup{poly}(\log V)\cdot\nu^{-1}\left(\lvert\langle\psi_{b}|\sqrt{\pi}\rangle\rvert^{-1}+\pi(E^{*})^{-\frac{1}{2}}e^{-\frac{b\eta}{2\alpha_{P}}}\right)\right).$
Hence if the short-jump can be performed in polynomial time, i.e.
$[\min(\textup{Gap}(-D(P)),\textup{Gap}(H_{b}))]^{-1}\lvert\langle\psi_{b}|\sqrt{\pi}\rangle\rvert^{-1}=\mathcal{O}\left(\text{poly}(n)\right),$
then the runtime of the overall generalized short path algorithm is $\mathcal{O}^{\ast}\left([\pi(E^{*})^{-1}]^{\left(\frac{1}{2}-\frac{b\eta}{2\alpha_{P}\ln(1/\pi(E^{*}))}\right)}\right)$.
@@ 5.1 Optimization with Fixed Hamming Weight: Transposition Walk
The $k$-particle Bernoulli-Laplace (BL) diffusion or Transposition Walk on $n$ sites is a random walk on the space of Hamming weight $k$ bitstrings. For our case we will work with $\pm 1$ strings or “spin configurations” $x$ and define the Hamming weight $\lvert x\rvert$ as the number of +1+1’s. Formally, $\mathcal{X}=\{x\in\{-1,1\}^{n}:\lvert x\rvert=k\}$. A single step of BL consists of choosing, uniformly at random, a transposition that swaps some $x_{j}=1$ with another $x_{i}=-1$.
@@
where $X_{j},Y_{j}$ denote the Pauli operators applied to qubit $j$. This is commonly called the complete-graph $XY$ mixer. Note that there is equality between $D(P)$ and $P$ because the walk is symmetric. The ground state of D(P)-D(P) is the uniform superposition over Hamming weight $k$ computational basis states, and thus encodes the stationary distribution of the transposition walk. Note that $|\sqrt{\pi}\rangle$ is just the Hamming weight $k$ Dicke state, which can be prepared efficiently [BE19].
@@
$\mathcal{D}(\psi,\psi)\geq\frac{n}{k(n-k)\tau_{0}\log\left(\frac{n}{\min(k,n-k)}\right)}\mathrm{Ent}(\psi^{2}),$
where $\tau_{0}$ is a universal constant.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the $k$-particle transposition walk on $n$ sites. Let $H:\left\{-1,1\right\}^{n}\rightarrow\mathbb{R}$ be a diagonal Hamiltonian with ground state energy $E^{\ast}$. If $\lVert H\rVert_{P}=\mathcal{O}(1)$, and
@@
then there exists a short path algorithm with runtime
@@
We apply the above result to a Hamming-weight constrained version of MaxCut over Erdős-Rényi graphs, which we call MaxCut-Hamming. One well-known special case is Hamming weight $\frac{n}{2}$ called MaxBisection.
5.1.1 Hamming weight Constrained MaxCut
Consider a graph $G(\mathcal{N},\mathcal{E})$ with vertex set $\mathcal{N}:=[n]$ and edge set $\mathcal{E}$. We assume $G$ is drawn from the Erdős-Rényi ensemble $\mathcal{G}\left(n,\frac{p}{n-1}\right)$ for a constant $p$, i.e., each edge $e_{ij}\in\mathcal{E}$ for $(i,j)\in[n]\times[n]$ is created with probability $\frac{p}{n-1}$ such that $G$ has an average degree of $p$. We are interested in solving the Maximum b
isection
problem:
@@ 5.1.1 Hamming weight Constrained MaxCut
which gives that $\frac{\lvert E^{*}\rvert}{\Delta_{P}}=\mathcal{O}(k)$. However, $\ln\binom{n}{k}=\Theta(k\ln(n/k))$ for $k=o(n)$. Assuming a tail bound, this leads to a speedup that is falling with $n$. Specifically, the speedup is falling with $\frac{1}{\ln(n)}$, which is reminiscent of the running time achieved by [Has18a]. We summarize the two cases in the following theorem.
@@
Let $\mathcal{M}=(\mathcal{X},P,\pi)$ be the $k$-particle transposition walk on $n$ sites. Let $H:\left\{-1,1\right\}^{n}\rightarrow\mathbb{R}$ be a diagonal Hamiltonian encoding the MaxCut-Hamming cost function for a graph $G\sim\mathcal{G}\left(n,\frac{p}{n-1}\right)$. Then either of the following runtimes hold depending on $k\leq n$.
  • If $k=\Theta(n)$, then there exists a short path algorithm with runtime
@@
  • if $k=o(n)$, then there exists a short path algorithm with runtime
@@ 5.2 Glauber Dynamics
Glauber dynamics is a Markov chain algorithm designed to sample from the Gibbs distribution of a system, particularly in spin models like the Ising or hardcore model [Gla63]. The Gibbs measure $\pi_{\beta}$ for a system with configuration space $\mathcal{X}$ and Hamiltonian $H$ is defined as
$\pi_{\beta}(x)=\frac{\exp(-\beta H(x))}{Z(\beta)}$
@@
Glauber dynamics generates a Markov chain with this measure as its stationary distribution by sequentially updating a single site (spin or vertex) according to marginal distribution. Specifically, the update proceeds as follows: (i) a site (vertex) $v$ is chosen uniformly at random; (ii) the states of all other sites $u\neq v$ remain unchanged; (iii) the new state of $v$ is sampled from the marginal distribution of $v$ conditioned on its neighbors.
@@
The efficiency of Gibbs sampling in this case is related to how fast the Glauber dynamics mixes to its stationary distribution. In fact, approximate Gibbs sampling is tightly connected to partition function estimation in terms of computational complexity [ŠVV09], both of which are closely related to statistical phase transitions. For bounded-degree graphs, these transitions have been shown correspond to the uniqueness/non-uniqueness threshold on an infinite $d$-regular tree which captures whether the root of the tree is affected by the leaves. Hence, this critical point is also called the tree-uniqueness threshold. In the uniqueness regime, correlations decay rapidly, allowing efficient approximation of the partition function. However, beyond the non-uniqueness threshold, long range correlations emerge and no polynomial-time algorithm can approximate the partition function. However, the tree threshold only appears to indicate the hardness transition for worst-case instances. The phase transition point for random graphs appears to be beyond the tree-uniqueness threshold.
@@
In particular, for the hardcore model with fugacity parameter $\lambda=e^{\beta}$, [Wei06] presented a fully polynomial time approximation scheme ($\mathsf{FPTAS}$) for computing the partition function for graphs with maximum degree $d$ when $\lambda\leq(1-\xi)\lambda_{c}$ where $\lambda_{c}=\frac{(d-1)^{(d-1)}}{(d-2)^{d}}$ is the corresponding uniqueness threshold. Additionally, Glauber dynamics is known to mix in polynomial-time for any constant $\xi$ and $d$. On the other hand, [Sly10] proved that there is no fully polynomial randomized approximation scheme ($\mathsf{FPRAS}$) to approximate the partition function for bounded-degree graphs when $\lambda>\lambda_{c}$ unless $\mathsf{NP}=\mathsf{RP}$ confirming the main conjecture of [MWW09].
For $\lambda$ at criticality, $\lambda=\lambda_{c}$, [CCYZ25] showed that Glauber dynamics mixes in polynomial time and hence partition function approximation is also polynomial time. However, approaching criticality $\lambda=(1-\xi)\lambda_{c}$, the modified log-Sobolev and standard log-Sobolev constants for Glauber dynamics appear to fall as $\mathcal{O}(\exp(-1/\xi))$ [CLV21]. Still, the Poincaré constant (hence the spectral gap) remains $\mathcal{O}(1/\text{poly}(n))$ at criticality.
For random regular graphs, the situation is quite different. [CCC+25] showed that Glauber dynamics mixes in polynomial time for $\lambda>\lambda_{c}$, going all the way up to $\lambda=\mathcal{O}(1/\sqrt{d})$. For the hardcore model, this was done by providing a Poincaré inequality rather than log-Sobolev inequality.
There is a similar situation for the antiferromagnetic Ising model over bounded-degree graphs. The critical inverse-temperature $\beta_{c}$, corresponding to the uniqueness/non-uniqueness phase transition on the infinite $d$-regular tree, is defined by $(d-1)\tanh(\beta_{c})=1$  [^6] Note in some literature the criticality condition is expressed in terms of $e^{-2\beta_{C}}$ where the condition becomes $e^{-2\beta_{C}}\geq\frac{d-2}{d}$. [CCYZ25]. [GSV16] show that unless $\mathsf{NP}=\mathsf{RP}$ there is no $\mathsf{FPRAS}$ for approximating the partition function of the antiferromagnetic Ising model when $\beta>\beta_{c}$. Glauber dynamics mixes in polynomial-time for $(d-1)\tanh(\beta_{c})\leq 1-\delta$ [CLV21] for constant $\delta\in(0,1)$. More recently, [CCYZ25] showed that Glauber dynamics also mixes efficiently for $(d-1)\tanh(\beta_{c})\leq 1+\frac{\alpha}{\sqrt{n}}$ for constant $\alpha>0$. Hence for the antiferromagnetic Ising model, we can mix slightly beyond the threshold with an edge that vanishes as $n\rightarrow\infty$.
For the antiferromagnetic Ising model over random $d$-regular graphs [CCC+25], Glauber dynamics mixes in total variation distance for constant $d$ when $(d-1)\tanh(\beta)\leq\frac{d-1}{8\sqrt{d-1}-1}$ [^7] This is converted from the condition $e^{-2\beta}\geq 1-\frac{1}{4\sqrt{d-1}}$ in [CCC+25, Corollary 5.5]. with high probability. This is close to a computational threshold, as it is known that with high probability no (stable [AMS24]) classical algorithm can efficiently mix in Wasserstein-2 distance beyond $\beta_{\textup{DC}}$, where $(d-1)\tanh(\beta_{\textup{DC}})=\frac{d-1}{\sqrt{d}}$ [HPP24], when $d$ is sufficiently large. This is due to a phenomenon known as disorder chaos [Cha09].
There is a well-known correspondence between the free-energy densities of the antiferromagnetic Ising model over random $d$-regular graphs, as $d\rightarrow\infty$, and the SK model [DMS17]. Specifically, the SK model also suffers from disorder chaos and no stable classical algorithm can efficiently mix in Wasserstein-2 when the inverse-temperature exceeds $\beta>1$ [AMS24, Theorem 2.5], with high probability. One can show that a properly normalized $\beta_{\textup{DC}}$ for the Ising case matches the SK point of disorder chaos, i.e. $\lim_{d\rightarrow\infty}\sqrt{d}\cdot\beta_{\textup{DC}}=1$. Classical stable algorithms have been shown to sample in Wasserstein-$2$ from the SK model for $\beta<1$ [AMS24, Cel22]. Our quantum algorithm provides a speedup for the regime of $\beta<\frac{1}{4}$ [CE22], where efficient sampling in TV can be shown.
@@
Let $P$ be a Glauber dynamics chain on $\mathcal{X}$ and let $H:\mathcal{X}\rightarrow\mathbb{R}$ be a diagonal Hamiltonian with ground state energy
@@
If the following holds for all $x,x^{\prime}$ such that $P(x,x^{\prime})>0$:
$\lvert H(x^{\prime})-H(x)\rvert\leq q,$ (38)
then $\|H\|_{P}\leq q^{2}$.
@@
$\|H\|_{P}$ $=\max_{x}\sum_{x^{\prime}}P(x,x^{\prime})(H(x^{\prime})-H(x))^{2}\leq\max_{x,x^{\prime}}\lvert H(x^{\prime})-H(x)\rvert^{2}\leq q^{2}.$
5.2.1 Bounding the Mixing time of Glauber Chains
In this section we review some of the known techniques for bounding the log-Sobolev and Poincaré constants of Glauber chains. These will be used to derive all of the results in the following subsections.
Consider the following definitions regarding an arbitrary measure $\pi$.
Definition 5.8 (Approximate Tensorization of Entropy).
We say that a distribution $\pi$ satisfies approximate tensorization of entropy with constant $C_{1}>0$ if for all $\pi$-measurable functions $f:\mathcal{X}\rightarrow\mathbb{R}$:
$\operatorname{Ent}_{\pi}[f]$ $\leq C_{1}\sum_{v\in\mathcal{X}}\operatorname{Ent}_{\pi_{v}}[f].$
Definition 5.9 ($u$-Marginally Bounded Measure (Definition 1.9 [CLV21])).
We say that the stationary distribution $\pi$ on $\mathcal{X}\subset\{-1,1\}^{n}$ is $u$-marginally bounded: if for every $\Lambda\subset\mathcal{X}$ and partial assigment $\tau:\Lambda\rightarrow\{-1,1\}$ that are not $\pi$-measure zero, it holds that for every $x\in\mathcal{X}\setminus\Lambda$:
$\pi(\sigma_{v}=i|\sigma_{\Lambda}=\tau)\geq u,$
whenever the above conditional measure is nonzero.
The above is used to provide alternative sufficient conditions to Theorem 4.17 from [CLV21] under which a modified log-Sobolev inequality can be promoted to a log-Sobolev inequality. We record this as the following fact.
Fact 5.10.
Let $\mathcal{X}$ be a set of size $n$ and µ be a distribution over $q]^{\mathcal{X}}$. If $\pi$ satisfies the approximate tensorization of entropy with constant $C_{1}$ and $\pi$ is $u$-marginally bounded, then the Glauber dynamics for $\pi$ satisfies the standard log-Sobolev inequality with constant $\omega=\frac{1-2u}{\log(1/u-1)}\frac{1}{C_{1}n}$ when $u<\frac{1}{2}$, or $\omega=\frac{1}{2C_{1}n}$ when $u=\frac{1}{2}$.
Proof.
Fix a configuration $x$ and consider a Markov chain $P_{v}$ that updates vertex $v$ according to marginal probability distribution $\pi_{v}=\pi(v|x_{\mathcal{X}-\{v\}})$. Then, LS constant of this Markov chain $\rho_{v}$ is lower bounded by $\rho_{v}\geq\frac{1-2\pi_{v}^{\ast}}{\log(1/\pi_{v}^{\ast}-1)}$ when $\pi_{v}^{\ast}<\frac{1}{2}$ or $\rho_{v}=\frac{1}{2}$ when $\pi_{v}^{\ast}=\frac{1}{2}$ due to [DSC96, Theorem A.1]. By the definition of the log-Sobolev constant, we have
The transition matrix of Glauber dynamics can be written as $\frac{1}{n}\sum_{v\in\mathcal{X}}\pi(v|\cdot)$. Therefore, the Dirichlet form for Glauber dynamics is
$\mathcal{D}(\sqrt{f},\sqrt{f})=\frac{1}{n}\sum_{v\in\mathcal{X}}\mathrm{Var}_{\pi_{v}}[\sqrt{f}].$
Using the tensorization of entropy,
$\operatorname{Ent}_{\pi}[f]$ $\leq C_{1}\sum_{v\in\mathcal{X}}\operatorname{Ent}_{\pi_{v}}[f]\leq C_{1}\sum_{v\in\mathcal{X}}\frac{1}{\rho_{v}}\mathrm{Var}_{\pi_{v}}[\sqrt{f}]\leq C_{1}n\max_{v\in\mathcal{X}}\left({\frac{1}{\rho_{v}}}\right)\mathcal{D}(\sqrt{f},\sqrt{f}).$
Combining the marginal boundedness property of $\pi$ (i.e., $\pi_{v}\geq u$ for all $v\in\mathcal{X}$), with the monotonicity of the function $\frac{1-2y}{\log(1/y-1)}$ for $y\in[0,1/2]$, it follows that $\omega\geq\frac{1-2u}{\log(1/u-1)}\frac{1}{C_{1}n}$ when $u<\frac{1}{2}$, and $\omega=\frac{1}{2C_{1}n}$ when $u=\frac{1}{2}$.
The approximate tensorization of entropy constant can be bounded if $\pi$ is marginally bounded and spectrally independent.
Definition 5.11 ($\rho$-Spectral Independence).
Given $\Lambda\subset V$ and partial assignment $\tau:\Lambda\rightarrow\{1,-1\}$ that is not $\pi$-measure zero. For every $(u,i),(v,j)\in V\times V$, $u\neq v$, if $\pi(\sigma_{w}=i|\sigma_{\Lambda}=\tau)>0,w\in\{u,v\}$, then we define the corresponding element of the influence matrix to be
$\Psi_{\pi}^{\tau}((u,i),(v,j))=\pi(\sigma_{v}=j|\sigma_{v}=i,\sigma_{\Lambda}=\tau)-\pi(\sigma_{v}=j|\sigma_{\Lambda}=\tau),$
and zero otherwise. We say that $\pi$ is $\eta$-spectrally independent if for every $\Lambda,\tau$, the largest eigenvalue of $\Psi_{\pi}^{\tau}$ is bounded by $\rho$.
Theorem 5.12 (Entropy Factorization, Theorem 2.9 in [CLV21]).
Let $d\geq 3$ be an integer and $b,\rho>0$ be reals. Suppose that $G=(\mathcal{N},\mathcal{E})$ is an n-vertex graph of maximum degree at most $d$ and $\mu$ is a totally connected Gibbs distribution of some spin system on $G$. If $\mu$ is both $u$-marginally bounded and $\rho$-spectrally independent and $n\geq\frac{24d}{u^{2}}(\frac{4\rho}{u^{2}}+1)$, then $\mu$ satisfies the approximate tensorization of entropy with constant
$C_{1}=\frac{18\log(1/u)}{u^{4}}\left(\frac{24d}{u^{2}}\right)^{\frac{4\rho}{u^{2}}+1}.$
Analogously to entropy, we have a similar notion of approximation tensorization of variance.
Definition 5.13 (Approximate Tensorization of Variance).
We say that a distribution $\pi$ satisfies approximate tensorization of variance with constant $C_{1}>0$ if for all $\pi$-measurable functions $f:\mathcal{X}\rightarrow\mathbb{R}$:
$\mathrm{Var}_{\pi}[f]$ $\leq C_{2}\sum_{v\in\mathcal{X}}\mathrm{Var}_{\pi_{v}}[f].$
Approximate tensorization of variance leads to a lower bound on the spectral gap of Glauber dynamics.
Fact 5.14 (Lemma 2.1 (1) from [CCYZ25]).
If $\pi$ satisfies $C_{2}$ approximate tensorization of variance, then the spectral gap of Glauber dynamics is lower bounded by $\frac{1}{C_{2}n}$.
5.2.2 Hardcore Model – Maximum Independent Set Problem